# Improved Supervised Fine-Tuning for Large Language Models to Mitigate Catastrophic Forgetting

### v1

@@ Abstract
Supervised Fine-Tuning (SFT), while enhancing large language models(LLMs)’ instruction-following capabilities and domain-specific task adaptability, often diminishes their general capabilities. Moreover, due to the inaccessibility of original pre-training data, catastrophic forgetting tends to be exacerbated when third-party practitioners implement SFT on open-sourced models. To address this challenge, we propose a novel, more cost-effective SFT method which could effectively reduce the risk of catastrophic forgetting without access to original SFT data. Our approach begins by reconstructing the likely SFT instruction distribution of the base model, followed by a multi-model screening process to select optimal data, which is then mixed with new data for SFT. Experimental results demonstrate that our method preserves generalization capabilities in general domains while improving task-specific performance.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable proficiency in understanding and generating human language, leading to their widespread adoption across diverse applications. Meanwhile, recent advancements have produced powerful, publicly accessible base LLMs, including LLaMA  Touvron et al. (2023a, b), Falcon Penedo et al. (2023), and Mistral  Jiang et al. (2023). To optimize performance in specialized domains such as medicine, law, and science, fine-tuning these general-purpose LLMs on domain-specific datasets is often necessary. However, this process introduces a critical challenge: catastrophic forgetting(McCloskey and Cohen, 1989), where models lose previously acquired general knowledge, thereby limiting their broader applicability and reusability across domains.
To mitigate catastrophic forgetting, a series of rehearsal-based methods have been proposed. However, traditional rehearsal-based methods mainly focus on the continual learning process and relies heavily on training data.To mitigate the degradation of general capabilities, the specific way is to blend general SFT data with domain-specific data, while mainstream open-source models like LlamA and Qwen do not disclose their original SFT datasets. Current practices often involve mixing public SFT datasets with custom data, but these public datasets frequently diverge significantly from the base model’s original SFT data distribution, exacerbating SFT performance degradation. Therefore, better SFT requires an alternative approach to deal with models that do not disclose their original SFT data. To address this challenge, we propose a novel, cost-effective rehearsal-based method that reconstructs the hidden instruction data distribution of base models and then generate synthetic SFT data through multi-response generation and filter for SFT. Furthermore, our method is compatible with existing continual learning approaches, enabling more effective sequential fine-tuning.
@@
  • We propose a method that could effectively reconstruct the likely hidden instruction distribution of open-source models by analyzing their generative behavior, guiding the creation of general-purpose SFT datasets to mitigate the degradation of general capabilities.
  • We introduce a Multi-Response Generation and Filter mechanism, which optimizes instruction-response pairs via mutual scoring across three models, significantly improving response quality and SFT efficacy.
  • Experiment results demonstrate that our SFT method can significantly reduce the decline in the LLM’s general capabilities after fine-tuning compared to directly mixing third-party datasets.
@@ 2 Related Work
2.1 Catastrophic Forgetting
Continual learning is essential for models that must adapt to dynamic environments, assimilating information from a continuous data stream while retaining previously learned knowledge. A critical obstacle in this domain is catastrophic forgetting, which refers to the pronounced degradation in performance on earlier tasks when the model is exposed to novel data. As the model adjusts its parameters to incorporate new information, it inadvertently overwrites previously acquired knowledge, thereby diminishing its performance on prior tasks. To address catastrophic forgetting, existing continual learning approaches can be broadly categorized into four classes: Replay-Based Shin et al. (2017); Ren et al. (2024), Regularization-Based Mi et al. (2020), Gradient-Based Lee et al. (2021), and Architecture-Based Geng et al. (2021) approaches.
@@ 3 Methods
Our framework comprises two stages: (1) reconstructing the original SFT data distribution from the base model, and (2) blending the reconstructed data with new domain-specific data for training.
3.1 Reconstructing General-Purpose SFT Data from LLMs
@@
The process of recovering the instruction distribution involves three automated steps: instruction generation, response generation, and high-quality response filtering.
Step 1: Instruction Generation. The objective of this step is to approximate the distribution of instructions used in the model’s original SFT phase. Given an open-weight aligned large language model (e.g., Llama-3-70B-Instruct), we design a pre-query template in the format of the predefined instruction template. We input the prompt "<|start_header_id|>user<|end_header_id|>" into the large model (Llama-3-70B-Instruct), which generates a single instruction in response. By repeating this process 100,000 times, we obtain a total of 100,000 instructions, which collectively represent the current instruction distribution of the large model.
Step 2: Multi-Response Generation. The objective of this step is to generate responses to the instructions obtained in Step 1.
We send these instructions to Llama-3-70B-Instruct and two additional powerful large language models (such as gpt4 and Qwen2.5-72B-Instruct). For each instruction, each model generates three responses, resulting in a total of nine responses for each instruction.
Step 3: Filtering High-quality Responses. For each instruction, we evaluate nine generated responses using the three models previously mentioned, assigning quality scores to each response. The scores from the three models are then averaged to identify the response with the highest overall score.
Combining the optimal response with the corresponding instruction forms the instruction dataset. The exact prompt we use for scoring is provided in Table 3.
3.2 Data mixing and training.
The reconstructed SFT data is blended with domain-specific SFT data at a controlled ratio, and the combined dataset is used for full fine-tuning.
@@ 4.1 Setup
We evaluate the performance of the general-purpose base LLM after domain-specific fine-tuning in the medical field. Our method-generated dataset is mixed with medical data for SFT, and performance is assessed on common general benchmarks. Comparisons are made against models fine-tuned using open-source third-party SFT datasets.
Base LLM Our experiment used Llama3-70B-Instruction as the base model. Llama3-70B-Instruction has undergone previous instruction fine-tuning during pre-train to improve its instruction following ability.
Baseline Datasets for SFT AWe adopted several common open-source SFT datasets instead of the SFT dataset established by our method mixed with the specific medical datasets as baselines. These datasets include: Evol Instruct (Xu et al., 2023), UltraChat (Ding et al., 2023), ShareGPT (Chiang et al., 2023), WildChat (Zhao et al., 2024),GenQA (Chen et al., 2024), OpenHermes 1 (Teknium, 2023b), OpenHermes 2.5 (Teknium, 2023a), and Tulu V2 Mix (Ivison et al., 2023). ShareGPT and WildChat are representative human-written datasets containing 112K and 652K high-quality multi-round conversations between humans and GPT, respectively. Evol Instruct, UltraChat, and GenQA are representative open-source synthetic datasets. Following Meng et al. (2024), we use the 208K sanitized version of Ultrachat provided by HuggingFace [^1] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k. OpenHermes 1, OpenHermes 2.5, and Tulu V2 Mix are crowd-sourced datasets consisting of a mix of diverse open-source instruction datasets, with 243K, 1M, and 326K conversations, respectively.
Evaluation Performance is measured on general-purpose tasks from the Hugging Face Open LLM Leaderboard (Beeching et al., 2023), including MMLU-PRO (general knowledge) (Wang et al., 2024), GPQA (expert-level Q&A) (Rein et al., 2023), IFEval (instruction following) (Zhou et al., 2023) and MATH level 5 (mathematical reasoning) (Hendrycks et al., 2021).
Implementation Details All datasets are standardized to the same total size, with medical data fixed at 17%. Training is conducted via LLaMA-Factory with a learning rate of 1e-6 and 1 epoch on 8 Tesla A800-80GB GPUs.
4.2 Results
As demonstrated in Table 1, the model fine-tuned with our method achieves marked improvements across all four general benchmarks compared to models trained on mixtures of third-party open-source datasets and domain-specific data.
Figure 1: SFT Data Distribution of Base LLM
Alignment Setup MMLU-PRO (5) GPQA (0) IFEval(0) Math Lvl 5 (4) Average
Llama-3-70B-Instruct 46.74 4.92 80.99 23.34 39.00
@@ 4.2 Results
Table 1: Final result of the SFT model using our method and the official instruction model across various downstream benchmarks.
Alignment Setup MMLU-PRO (5) GPQA (0) IFEval(0) Math Lvl 5 (4) Average
Llama-3-70B-Instruct 46.74 4.92 80.99 23.34 39.00
Single model-Single response 46.11 4.72 81.31 23.11 38.81
Single model-Multi-responses-Filtered 46.55 4.91 81.72 23.06 39.06
Three Models-Multi-responses-Filtered 46.73 4.88 81.93 23.29 39.21
Table 2: Ablation result of our method.
Impact of SFT Data Distribution As illustrated in Figure 1, the distribution of SFT instruction types extracted from the original LLM primarily comprises information-seeking and advice-seeking data, with part proportions of coding data, a small amount of mathematical and reasoning tasks, and other categories. Our reconstructed general dataset, directly generated from the base model’s instructions, strictly adheres to this intrinsic distribution. In contrast, existing open-source SFT datasets exhibit more significant divergence from the original distribution, particularly in their underrepresentation of coding and reasoning tasks, and overrepresentation of generic conversational templates. This misalignment exacerbates catastrophic forgetting when third-party datasets are used for fine-tuning, as the model’s parameters are forced to adapt to a distorted data landscape that conflicts with its pre-training and initial SFT phases.
@@ 4.3 Ablation Study
To further examine the impact of individual components in our proposed method, we conduct an ablation study focusing on multi-response generation and filtering high-quality responses.We respectively test another two configurations:
Single model-Single response After instruction generation, only the target model (LlamA3-70B-Instruct) generates a single response, bypassing multi-response generation and filtering.
Single model-Multi-responses-Filtered The target model generates three responses without external models, followed by filtering.
Results are presented in table 2. The single-response setup exhibits performance degradation compared to the full method but still surpasses models fine-tuned with open-source dataset mixtures in table 1 with the three-response setup (all generated by one LLM without external models and filtered) albeit slightly underperforming the complete method.
These findings highlight the critical role of all three stages in our framework. Instruction Generation serves as the foundation, ensuring the data distribution approximates the original. Subsequently, Multi-Response Generation and Filtering High-Quality Responses iteratively refine data quality by synthesizing multiple candidate responses and selecting optimal outputs. Crucially, integrating more capable LLMs (GPT-4 and Qwen2.5-72B-Instruct) as response generators introduces a knowledge distillation effect, where the superior reasoning and generation capabilities of these models enhance the quality of synthesized SFT data. This hierarchical design—first recovering the latent instruction distribution, then refining responses through multi-model collaboration—ensures both distributional fidelity and high-quality supervision signals, which are essential for mitigating catastrophic forgetting while improving task-specific performance.
@@ 5 Conclusion
In this paper, we developed a method to extract instruction distributions from a model trained on an unpublished instruction dataset. We then leveraged two additional powerful models to collaboratively generate high-quality responses, forming a dataset used for SFT. Compared to other baseline open-source SFT datasets, our approach mitigates more catastrophic forgetting and enhances general fine-tuning performance.
6 Declaration
@@
Appendix A A prompt used to evaluate the quality of a response.

### v2

@@ Abstract
Supervised Fine-Tuning (SFT) is a critical step for enhancing the instruction-following capabilities of Large Language Models (LLMs) and adapting them to specialized domains. However, SFT often leads to a degradation of the model’s general abilities, a phenomenon known as catastrophic forgetting. This problem is exacerbated when third-party practitioners fine-tune open-source models, as the original SFT data is typically not available. To address this challenge, we propose a novel and cost-effective SFT method that effectively mitigates catastrophic forgetting without requiring access to the original SFT data. Our approach first reconstructs the likely instruction distribution of the base model. It then employs a multi-model generation and filtering pipeline to synthesize a high-quality general-purpose dataset. This synthetic dataset is mixed with new, domain-specific data for fine-tuning. Experimental results show that our method not only preserves the model’s capabilities in general domains but also improves task-specific performance, outperforming baselines that use publicly available SFT datasets.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable proficiency in understanding and generating human-like text, leading to their widespread adoption across diverse applications. The release of powerful, publicly accessible base models like LLaMA (Touvron et al., 2023a, b), Falcon (Penedo et al., 2023), and Mistral (Jiang et al., 2023) has catalyzed a wave of specialization. To optimize these models for specific domains such as medicine, law, or finance, practitioners often perform Supervised Fine-Tuning (SFT) on domain-specific datasets.
However, this specialization introduces a critical challenge: catastrophic forgetting (McCloskey and Cohen, 1989). As the model’s parameters adapt to the new data distribution, it tends to lose previously acquired general knowledge, thereby diminishing its performance on a wide range of tasks and limiting its broader applicability.
A common strategy to mitigate catastrophic forgetting is to use rehearsal-based methods, which involve mixing general-purpose data with domain-specific data during fine-tuning. However, the effectiveness of this approach hinges on the quality and distributional similarity of the rehearsal data to the data used for the model’s initial alignment. Mainstream open-source models do not disclose their original SFT datasets, forcing practitioners to rely on public SFT datasets. These datasets often diverge significantly from the base model’s original SFT data distribution, which can lead to suboptimal performance and even exacerbate knowledge degradation.
To overcome this limitation, we propose a novel, cost-effective rehearsal-based method designed for scenarios where the original SFT data is inaccessible. Our approach comprises two main stages. First, we reconstruct the implicit instruction data distribution of the base model. Second, we leverage this distribution to guide the synthesis of a high-quality, general-purpose SFT dataset using a multi-model generation and filtering process. This synthetic dataset serves as a proxy for the original alignment data and is then mixed with new domain-specific data for fine-tuning. Our method is also compatible with existing continual learning frameworks, offering a robust solution for sequential fine-tuning.
@@
  • We propose a novel method to reconstruct the latent instruction distribution of an aligned LLM by analyzing its generative behavior, enabling the creation of a general-purpose SFT dataset that mitigates catastrophic forgetting.
  • We introduce a Multi-Response Generation and Filter pipeline that uses a committee of powerful LLMs to generate and score responses, significantly improving the quality and diversity of the synthetic SFT data.
  • Experimental results demonstrate that our method significantly reduces the decline in an LLM’s general capabilities after domain-specific fine-tuning compared to standard approaches that mix in third-party datasets.
@@ 2 Related Work
2.1 Catastrophic Forgetting in LLMs
Continual learning aims to enable models to learn from a continuous stream of data without forgetting previously acquired knowledge. A primary obstacle in this field is catastrophic forgetting, the sharp decline in performance on past tasks after a model is trained on new data. As the model updates its parameters to fit the new data, it can overwrite the weights crucial for performing earlier tasks. In the context of LLMs, this manifests as a loss of general knowledge and reasoning skills after fine-tuning on a narrow, specialized dataset.
To combat catastrophic forgetting, researchers have proposed several categories of methods: Replay-Based (Shin et al., 2017; Ren et al., 2024), Regularization-Based (Mi et al., 2020), Gradient-Based (Lee et al., 2021), and Architecture-Based (Geng et al., 2021) approaches. Our work falls under the replay-based paradigm but innovates by synthesizing the replay data directly from the model to be trained.
@@ 3 Methods
Our proposed framework aims to create a high-quality, general-purpose dataset that mimics the base model’s original alignment data. This process consists of two main stages: (1) reconstructing the general-purpose SFT data from the base LLM and (2) mixing this reconstructed data with new domain-specific data for fine-tuning.
3.1 Reconstructing General-Purpose SFT Data
@@
The reconstruction process involves three automated steps: instruction generation, multi-response generation, and high-quality response filtering.
3.1.1 Step 1: Instruction Generation
The goal of this step is to produce a set of instructions $\{I_{1},I_{2},\dots,I_{N}\}$ whose distribution approximates that of the instructions used in the base model’s original SFT phase. We hypothesize that an aligned model’s generative priors are shaped by its training data. Therefore, by prompting the model to generate instructions, we can sample from a distribution that is implicitly conditioned by its alignment.
Given an open-weight aligned LLM, $M_{\text{base}}$ (e.g., Llama-3-70B-Instruct), we use a simple pre-query template to elicit instruction generation. Specifically, we feed the model the prompt corresponding to the user’s turn in its conversational template, such as "<|start_header_id|>user<|end_header_id|>" for Llama-3. The model’s continuation is treated as a generated instruction. We repeat this process $N$ times to obtain a large corpus of instructions, $D_{\text{instr}}=\{I_{1},\dots,I_{N}\}$. In our experiments, we set $N=100,\!000$ to ensure a diverse and representative sample.
3.1.2 Step 2: Multi-Response Generation
Once we have the set of instructions $D_{instr}$, the next step is to generate high-quality responses for them. A single model, even the base model itself, may produce suboptimal or generic responses. To enhance quality and diversity, we employ a multi-model generation strategy.
For each instruction $I_{j}\in D_{instr}$, we query a committee of $K$ powerful LLMs, $\{M_{1},M_{2},\dots,M_{K}\}$, to generate responses. In our setup, we use the base model $M_{base}$ itself, along with two other state-of-the-art models known for their strong performance (e.g., GPT-4 and Qwen2.5-72B-Instruct). For each instruction $I_{j}$, each model $M_{k}$ generates $L$ candidate responses $\{R_{j,k,1},\dots,R_{j,k,L}\}$ using nucleus sampling to encourage diversity. This results in a total of $K\times L$ candidate responses for each instruction. We use $L=3$, yielding nine responses per instruction.
3.1.3 Step 3: Filtering High-Quality Responses
With a pool of candidate responses for each instruction, the final step is to select the best one. We leverage the same committee of LLMs, $\{M_{1},\dots,M_{K}\}$, as judges. For each instruction $I_{j}$, we task each model $M_{k}$ with evaluating all $K\times L$ candidate responses.
The evaluation is guided by a carefully designed prompt, shown in Table 3 in the Appendix. This prompt asks the model to score a response on a 5-point scale based on criteria such as helpfulness, relevance, clarity, and adherence to an AI persona. For each candidate response $R_{j,k^{\prime},l}$, we obtain a set of scores $\{s_{j,k^{\prime},l,1},\dots,s_{j,k^{\prime},l,K}\}$ from the judging models. The final score for the response is the average:
$\bar{s}_{j,k^{\prime},l}=\frac{1}{K}\sum_{k=1}^{K}s_{j,k^{\prime},l,k}$
We then select the response with the highest average score as the optimal response, $R_{j}^{*}$:
$R_{j}^{*}=\underset{k^{\prime}\in\{1,\dots,K\},l\in\{1,\dots,L\}}{\text{argmax}}\bar{s}_{j,k^{\prime},l}$
The final reconstructed general-purpose dataset is composed of the instruction-response pairs: $D_{recon}=\{(I_{1},R_{1}^{*}),\dots,(I_{N},R_{N}^{*})\}$.
3.2 Data Mixing and Training
The reconstructed general-purpose dataset, $D_{recon}$, serves as a high-quality proxy for the original SFT data. To adapt the model to a new domain, we mix $D_{recon}$ with the new domain-specific SFT dataset, $D_{domain}$. The final training dataset is $D_{train}=D_{recon}\cup D_{domain}$. The mixing ratio can be tuned as a hyperparameter, but we found that even a simple concatenation works well. We then perform full fine-tuning of the base model on $D_{train}$.
@@ 4.1 Setup
Our experiments are designed to evaluate whether fine-tuning with our reconstructed dataset can effectively mitigate catastrophic forgetting in a domain-specialization task. We assess performance on a suite of general-knowledge benchmarks after fine-tuning on medical data.
Base LLM. We use Llama-3-70B-Instruct as our base model. This model has been instruction-tuned by its developers and exhibits strong general capabilities, making it a suitable candidate for studying the effects of catastrophic forgetting.
Domain-Specific Data. We use a proprietary medical question-answering dataset as our target domain data, $D_{domain}$.
Baseline Datasets for Rehearsal. To establish strong baselines, we compare our method against the standard practice of mixing publicly available SFT datasets with the domain-specific data. We selected a diverse set of popular open-source datasets: ShareGPT (Chiang et al., 2023) and WildChat (Zhao et al., 2024) (human-written); Evol-Instruct (Xu et al., 2023), UltraChat (Ding et al., 2023), and GenQA (Chen et al., 2024) (synthetic); OpenHermes 1 (Teknium, 2023b), OpenHermes 2.5 (Teknium, 2023a), and Tulu V2 Mix (Ivison et al., 2023) (mixed-source). For fair comparison, we follow Meng et al. (2024) and use the 208K sanitized version of UltraChat.
Evaluation. We measure the model’s performance on a broad range of general-purpose tasks from the Hugging Face Open LLM Leaderboard (Beeching et al., 2023). The selected benchmarks test different capabilities:
  • MMLU-PRO (Wang et al., 2024): Advanced reasoning and general knowledge.
  • GPQA (Rein et al., 2023): Graduate-level and expert-level question answering.
  • IFEval (Zhou et al., 2023): Instruction following capability.
  • MATH (Level 5) (Hendrycks et al., 2021): Complex mathematical reasoning.
Implementation Details. For all experiments, we create training datasets of the same total size. The domain-specific medical data constitutes a fixed 17% of the final mix, with the remaining 83% being the general-purpose rehearsal data (either from our method or a baseline dataset). We perform full fine-tuning for 1 epoch using the LLaMA-Factory library. The training was conducted on 8 NVIDIA A800-80GB GPUs. We used the AdamW optimizer with a learning rate of $1\times 10^{-6}$, a global batch size of 64, and a sequence length of 4096.
Table 1: Main results comparing our method against the baselines using public SFT datasets for rehearsal. All models were fine-tuned on a mix of general and medical data . Our method achieves the highest average score, demonstrating its effectiveness in mitigating catastrophic forgetting.
Alignment Setup MMLU-PRO GPQA IFEval MATH Lvl 5 Average
Llama-3-70B-Instruct (Base Model) 46.74 4.92 80.99 23.34 39.00
@@
4.2 Results
The main experimental results are presented in Table 1. Our method, which uses the reconstructed dataset for rehearsal, not only prevents a decline in general capabilities but also achieves a slight improvement over the original Llama-3-70B-Instruct model.
As shown in the table, all baseline models that were fine-tuned using public datasets exhibit a drop in average performance compared to the original Llama-3-70B-Instruct model. The degradation is particularly severe for math and reasoning-heavy tasks like MATH and GPQA. For instance, using GenQA results in a drop of over 3 points in the average score. This confirms our hypothesis that fine-tuning with distributionally mismatched data exacerbates catastrophic forgetting.
In stark contrast, the model fine-tuned with our reconstructed dataset achieves an average score of 39.21, surpassing the original model’s score of 39.00. Our method nearly perfectly preserves the challenging MMLU-PRO and MATH scores while maintaining strong performance on GPQA and IFEval. This demonstrates the critical importance of using rehearsal data that aligns with the model’s original instruction distribution.
Impact of SFT Data Distribution. To understand why our method works, we analyzed the distribution of instruction types in our reconstructed dataset. As illustrated in Figure 1, the data recovered from Llama-3-70B-Instruct contains a significant proportion of coding, mathematical, and reasoning tasks, alongside more common information-seeking queries. This reflects the diverse and challenging data likely used in its original alignment. In contrast, many public SFT datasets over-represent generic conversational queries and under-represent complex reasoning tasks. This distributional mismatch forces the model’s parameters to shift away from the regions that encode these critical skills, leading to performance degradation. Our method succeeds by first recovering a faithful data distribution and then populating it with high-quality content.
Figure 1: Task category distribution of the SFT data reconstructed from the base LLM (Llama-3-70B-Instruct).
Table 2: Ablation study results. The full method outperforms the simplified versions, highlighting the benefits of both multi-response generation and using a diverse committee of models for generation and filtering.
Alignment Setup MMLU-PRO GPQA IFEval MATH Lvl 5 Average
Llama-3-70B-Instruct (Base Model) 46.74 4.92 80.99 23.34 39.00
Single-Model-Single-Response 46.11 4.72 81.31 23.11 38.81
Single-Model-Multi-Response-Filtered 46.55 4.91 81.72 23.06 39.06
Ours (Three-Models-Multi-Response-Filtered) 46.73 4.88 81.93 23.29 39.21
@@ 4.3 Ablation Study
To dissect the contribution of each component in our method, we conducted an ablation study focusing on the multi-response generation and filtering stages. We tested two simplified configurations:
  • Single-Model-Single-Response: Instructions are generated as before, but responses are generated only by the base model ($M_{base}$), with one response per instruction and no filtering.
  • Single-Model-Multi-Response-Filtered: Responses are generated only by the base model ($L=3$), which then also acts as the sole judge to filter for the best response.
The results are presented in Table 2. The Single-Model-Single-Response setup, while being the simplest, already achieves an average score of 38.81. This is a strong result, outperforming nearly all public dataset baselines in Table 1 and underscoring the fundamental importance of getting the instruction distribution right.
Moving to the Single-Model-Multi-Response-Filtered setup, we see a notable improvement to an average score of 39.06. This demonstrates the value of generating multiple candidate responses and applying a quality filter, even when using only the base model. This step helps to mitigate instances where the model might produce a low-quality or incomplete initial response.
Finally, our full method, Three-Models-Multi-Response-Filtered, achieves the best performance at 39.21. The improvement from 39.06 to 39.21 highlights the benefit of using a committee of diverse and powerful models. By incorporating stronger models like GPT-4, we are effectively performing a form of knowledge distillation. These superior models provide higher-quality responses, which serve as better training signals, and also act as more discerning judges, further elevating the quality of the final reconstructed dataset. This hierarchical refinement—first matching the distribution, then enhancing quality through multi-model collaboration—is key to our method’s success.
@@ 5 Conclusion
In this paper, we introduced a novel method to mitigate catastrophic forgetting during supervised fine-tuning when the original alignment data is unavailable. Our approach reconstructs the base model’s latent instruction distribution and then synthesizes a high-quality, general-purpose dataset using a multi-model generation and filtering pipeline. When mixed with domain-specific data, this reconstructed dataset allows the LLM to specialize without sacrificing its general capabilities. Our experiments show that this approach is significantly more effective than the common practice of using public SFT datasets for rehearsal. The model fine-tuned with our method not only avoids performance degradation but even surpasses the original base model’s average score on a challenging set of general benchmarks.
@@
Declaration
Appendix A Prompt for Response Quality Evaluation
The prompt used to have the committee of LLMs score the quality of a generated response is detailed in Table 3.
Table 3: The prompt used to evaluate the quality of a response.